OpenStudy (anonymous):

(3a+2b)(3a-2b)(9a^2+4b^2) ?

OpenStudy (anonymous):

Do you want it simplified or expanded?

OpenStudy (mathmale):

Note that (a+b)(a-b)=a^2 - b^2. Use this fact to multiply the first two factors together.

OpenStudy (anonymous):

Simplified @larryboxaplenty thanks @mathmale

OpenStudy (anonymous):

I meant to ask if you need it in expanded form or root form

OpenStudy (anonymous):

E.g. like 3x^2+2x+1 or (x+1)(x-1)(x+10)(x-4)

OpenStudy (anonymous):

First one @larryboxaplenty

OpenStudy (anonymous):

first do a (3a-2b)(9a^2+4b^2)

OpenStudy (anonymous):

It's easier to do 1-term * 2 term first than 2 term * 2 term

OpenStudy (anonymous):

Ok thanks! @larryboxaplenty

OpenStudy (mathmale):

Alternative approach: (3a+2b)(3a-2b)(9a^2+4b^2) = (9a^2 - 4b^2)*(9a^2 + 4b^2) = 81a^4 - 16b^2. That's it.

OpenStudy (anonymous):

Thanks! So easy to understand your explanation :) @mathmale